# QWENDY: Gene Regulatory Network Inference Enhanced by Large Language Model and Transformer

### v1

@@ Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. It is based on a gene expression model and solves the transformation for the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. Then we enhance QWENDY by transformer neural networks and large language models to obtain two variants: TEQWENDY and LEQWENDY. We test the performance of these methods on two experimental data sets and two synthetic data sets. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets.
@@ 1 Introduction
One gene can activate or inhibit the expression of another gene. Genes and such regulation relations form a gene regulatory network (GRN). For $n$ genes, the corresponding GRN is commonly expressed as an $n\times n$ matrix $A$, where $A_{ij}>0/=0/<0$ means gene $i$ has positive/no/negative regulation effect on gene $j$. If the regulation relations of concerned genes are known, one can understand the corresponding biological process or even control it with gene perturbation. Therefore, knowledge of GRNs can be useful to developmental biology [1, 2, 3], and even studying macroscopic behavior [4, 5, 6]. Since it is difficult to determine the GRN directly, the common practice is to infer the GRN from gene expression data.
With new measurement techniques, such as scRNA-seq, one can measure the expression levels (mRNA counts) of different genes for a single cell. Since gene expression at the single cell level is random, this measurement can be repeated for different cells, and we can use those samples to obtain the probability distribution of different genes at this time point. Since the measurement kills cells, one cell can only be measured once. Therefore, we cannot obtain the joint probability distribution of gene expression levels at different time points. When we measure at multiple time points, we can only obtain a marginal probability distribution for each time point.
@@
WENDY method is derived from a simplified gene expression model, and the performance of WENDY is not satisfactory when the model does not fit with experiments. Instead of developing a more complicated gene expression model, we developed TRENDY method [10], which trains a transformer, a deep learning architecture that can be applied to various fields [11], to transform the input data to better fit the gene expression model. Then the transformed data will produce more accurate GRNs by WENDY. Such more accurate GRNs will be further enhanced by another transformer model.
@@
We test the performance of QWENDY, TEQWENDY, and LEQWENDY on two experimental data sets and two synthetic data sets. The TRENDY paper [10] has tested 16 methods on the same data sets, meaning that we can compare them with our three new methods. In all 19 GRN inference methods, TEQWENDY has the best overall performance . If we restrict to two experimental data sets, QWENDY ranks the first. Therefore, we recommend applying QWENDY or TEQWENDY.
@@ 2 Literature review
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [14] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [15, 16] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [17] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, we only know one model-based method, WENDY.
@@ 3.2 Algorithm details
$L_{1}L_{1}^{\mathrm{T}}-L_{1}OO^{\mathrm{T}}L_{1}^{\mathrm{T}}.$
@@
$\begin{split}&||D_{2}-WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}=||D_{2}||_{\mathrm{F}}^{2}+||WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}[D_{2}\otimes(WD_{1}W^{\mathrm{T}})]_{ii}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}d_{2,i}\sum_{j=1}^{n}W_{ij}^{2}d_{1,j}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}}.\end{split}$
Thus we just need to maximize $\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}}$. Notice that $W\otimes W$ is doubly-stochastic, meaning that it is non-negative, and each row or column has sum $1$. By Birkhoff–von Neumann Theorem, $W\otimes W$ can be decomposed to
@@
Due to the rearrangement inequality, for each permutation matrix $Q_{i}$,
@@
$\begin{split}&||P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2}-WP_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1}W||_{\mathrm{F}}^{2}\\ =&||G-WHW||_{\mathrm{F}}^{2}=||G||_{\mathrm{F}}^{2}+||WHW||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}\\ =&||G||_{\mathrm{F}}^{2}+||H||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2},\end{split}$ (11)
where $\otimes$ is element-wise product. We want to find diagonal $W$ with ±$\pm 1$ that minimizes Eq. 10, which is equivalent to maximizing $||G\otimes(WHW)||_{\mathrm{F}}^{2}$. Define $C=G\otimes H$, which is still positive definite and symmetric by Schur product theorem. Assume that $C$ does not have repeated eigenvalues. Denote the diagonal elements of $W$ by $\mathbf{w}=[w_{1},\ldots,w_{n}]$. Then we need to maximize
@@
Step (4): For $B$ and $-B$ from Step (3), we determine which satisfies Eqs. 46 better. Notice that Eqs. 46 share an unknown $\mathbf{c}$. Define
@@ 3.3 Correctness of QWENDY method
$D_{1}$ has the same eigenvalues as
$L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}=L_{0}^{-1}B_{0}^{\mathrm{T}}L_{0}L_{0}^{\mathrm{T}}B_{0}L_{0}^{-\mathrm{T}}.$
Since $L_{0}^{-1}B_{0}^{\mathrm{T}}L_{0}$ has the same eigenvalues of $B_{0}^{T}$, $D_{1}$ has the same eigenvalues of $B_{0}^{T}B_{0}$. The same argument holds for $D_{2}$. Therefore, $D_{1}$ and $D_{2}$ (with increasing diagonal values) are equal. Assume that the diagonal elements $d_{1},\ldots,d_{n}$ of $D_{1}$ are distinct. Define $W_{0}=P_{2}^{\mathrm{T}}O_{0}P_{1}$, then
@@
and $W_{0}[i,j]d_{i}=W_{0}[i,j]d_{j}$. For $i\neq j$, $d_{i}\neq d_{j}$, meaning that $W_{0}[i,j]=0$. Now $W_{0}$ is diagonal and orthonormal, implying that its diagonal elements are $1$ or $-1$, and
@@
namely the eigenvector of the largest eigenvalue. The projection $w_{i}=\mathrm{sign}(v_{i})$ has no effect, and
@@ 4 Enhanced QWENDY method
For this problem, TRENDY method proposes that we can construct $K_{1}^{*}=B^{\mathrm{T}}K_{0}B$, and train a model with input $K_{1}$ and target $K_{1}^{*}$, so that the output $K_{1}^{\prime}$ is close to $K_{1}^{*}$. Then we can apply WENDY to $K_{0}$, $K_{1}^{\prime}$ to obtain a more accurate GRN $A_{1}$. After that, we can train another model with input $K_{0}$, $K_{1}$, $A_{1}$, and the target is the true GRN $A_{\text{true}}$. Then the final output $A_{2}$ is more similar to the true GRN than $A_{1}$.
@@
    Repeat generating random $A_{\mathrm{true}}$ and corresponding gene expression data at time 00, $t$, $2t$, $3t$
@@
    Input: gene expression data at four time points
@@ 5 Performance of different methods
The same as in the WENDY paper and the TRENDY paper [9, 10], we test different methods on two synthetic data sets and two experimental data sets. SINC data set [10] and DREAM4 data set [41] are generated by simulating a stochastic differential equation system. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [42]. hESC data set is from human embryonic stem cell-derived progenitor cells[43]. THP-1 and hESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
@@
We test the performance of QWENDY, TEQWENDY, and LEQWENDY methods on these four data sets. Since TRENDY ranks the first among 16 methods on these four data sets [10], we also present the performance of TRENDY. For each method, we compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores. These two scores, both between 00 and $1$, evaluate the level of matching, where $1$ means perfect match, and 00 means perfect mismatch.
@@ 6 Discussion
In this paper, we present QWENDY, a GRN inference method that requires single-cell gene expression data measured at four time points. Then we train the transformer-based variant TEQWENDY from scratch and fine-tune the LLM-based variant LEQWENDY based on pre-trained weights. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets. Notice that each experimental data set only has one group of data, meaning that the performance of each method has a large uncertain level.
@@ Acknowledgments
For all models in TEQWENDY and LEQWENDY, the loss function is mean squared error; the optimizer is Adam with learning rate $0.001$; the number of training epochs is $100$; the training stops early and rolls back to the best status if there is no improvement for consecutive $10$ epochs.

### v2

@@ Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. Based on a linear gene expression model, it solves the transformation for the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. Then we enhance QWENDY by the transformer neural networks and large language models to obtain two variants: TEQWENDY and LEQWENDY. We test the performance of these methods on two experimental data sets and two synthetic data sets. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets.
@@ 1 Introduction
One gene can activate or inhibit the expression of another gene. Genes and such regulation relations form a gene regulatory network (GRN). For $n$ genes, the corresponding GRN is commonly expressed as an $n\times n$ matrix $A$, where $A_{ij}>0/=0/<0$ means gene $i$ has positive/no/negative regulation effect on gene $j$, and its absolute value represents the regulation strength. If the regulation relations of concerned genes are known, one can understand the corresponding biological process or even control it with gene perturbation. Therefore, knowledge of GRNs can be useful in developmental biology [1, 2, 3], and even in the study of macroscopic behavior [4, 5, 6]. Since it is difficult to determine the GRN directly, the common practice is to infer the GRN from gene expression data.
With new measurement techniques, such as scRNA-seq, one can measure the expression levels (mRNA counts) of different genes for a single cell. Since gene expression at the single cell level is random, this measurement can be repeated for different cells, and we can use those samples to obtain the probability distribution of different genes at this time point. Since the measurement kills cells, one cell can only be measured once. Therefore, we cannot obtain the joint probability distribution of gene expression levels at different time points. When we measure at multiple time points, we can only obtain a marginal probability distribution of different genes for each time point.
@@
WENDY method is derived from a simplified gene expression model, and the performance of WENDY is not satisfactory when the model does not fit with experiments. Instead of developing a more complicated gene expression model, we developed TRENDY method [10], which trains a transformer model, a deep learning architecture that can be applied to various fields [11], to transform the input data to better fit the gene expression model. Then the transformed data will produce more accurate GRNs by WENDY. Such more accurate GRNs will be further enhanced by another transformer model.
@@
We test the performance of QWENDY, TEQWENDY, and LEQWENDY on two experimental data sets and two synthetic data sets. The TRENDY paper [10] has tested 16 methods on the same data sets, meaning that we can compare them with our three new methods. In all 19 GRN inference methods, TEQWENDY has the best overall performance. If we restrict ourselves to two experimental data sets, QWENDY ranks first. Therefore, we recommend applying QWENDY or TEQWENDY.
@@ 2 Literature review
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [14] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [15, 16] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [17] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, where each cell is measured only once, we only know one model-based method, WENDY.
@@ 3.2 Algorithm details
$||L_{1}L_{1}^{\mathrm{T}}-L_{1}OO^{\mathrm{T}}L_{1}^{\mathrm{T}}||_{\mathrm{F}}^{2}.$
@@
$\begin{split}&||D_{2}-WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}=||D_{2}||_{\mathrm{F}}^{2}+||WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}[D_{2}\otimes(WD_{1}W^{\mathrm{T}})]_{ii}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}d_{2,i}\sum_{j=1}^{n}W_{ij}^{2}d_{1,j}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}},\end{split}$
where $\otimes$ is the element-wise product. Thus we just need to maximize $\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}}$. Notice that $W\otimes W$ is doubly-stochastic, meaning that it is non-negative, and each row or column has sum $1$. By Birkhoff–von Neumann Theorem [37], $W\otimes W$ can be decomposed to
@@
Due to the rearrangement inequality [38], for each permutation matrix $Q_{i}$,
@@
$\begin{split}&||P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2}-WP_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1}W||_{\mathrm{F}}^{2}\\ =&||G-WHW||_{\mathrm{F}}^{2}=||G||_{\mathrm{F}}^{2}+||WHW||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}\\ =&||G||_{\mathrm{F}}^{2}+||H||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}.\end{split}$ (11)
We want to find diagonal $W$ with ±$\pm 1$ that minimizes Eq. 10, which is equivalent to maximizing $||G\otimes(WHW)||_{\mathrm{F}}^{2}$. Define $C=G\otimes H$, which is still positive definite and symmetric by Schur product theorem [39]. Assume that $C$ does not have repeated eigenvalues. Denote the diagonal elements of $W$ by $\mathbf{w}=[w_{1},\ldots,w_{n}]$. Then we need to maximize
@@
Step (4): For $B$ and $-B$ from Step (3), we determine which satisfies Eqs. 46 better. Notice that Eqs. 46 share the same unknown $\mathbf{c}$. Define
@@ 3.3 Correctness of QWENDY method
Since $P_{2}^{\mathrm{T}}O_{0}P_{1}$ is orthonormal, $D_{1}$ and $D_{2}$ are similar and have the same eigenvalues. Therefore, $D_{1}$ and $D_{2}$ as diagonal matrices with increasing diagonal values are equal. Assume that the diagonal elements $d_{1},\ldots,d_{n}$ of $D_{1}$ are distinct. Define $W_{0}=P_{2}^{\mathrm{T}}O_{0}P_{1}$, then
@@
and $W_{0}[i,j]d_{i}=W_{0}[i,j]d_{j}$. For $i\neq j$, we have $d_{i}\neq d_{j}$, meaning that $W_{0}[i,j]=0$. Now $W_{0}$ is diagonal and orthonormal, implying that its diagonal elements are $1$ or $-1$, and
@@
namely the eigenvector of the largest eigenvalue. Since $w_{i}=\pm 1$, the projection $w_{i}=\mathrm{sign}(v_{i})$ has no effect, and
@@ 4 Enhanced QWENDY method
For this problem, TRENDY method [10] proposes that we can construct $K_{1}^{*}=B^{\mathrm{T}}K_{0}B$, and train a model with input $K_{1}$ and target $K_{1}^{*}$, so that the output $K_{1}^{\prime}$ is close to $K_{1}^{*}$. Then we can apply WENDY to $K_{0}$, $K_{1}^{\prime}$ to obtain a more accurate GRN $A_{1}$. After that, we can train another model with input $K_{0}$, $K_{1}$, $A_{1}$, and the target is the true GRN $A_{\text{true}}$. Then the final output $A_{2}$ is more similar to the true GRN than $A_{1}$.
@@
    Repeat generating random GRN $A_{\mathrm{true}}$ and corresponding gene expression data at time 00, $t$, $2t$, $3t$ from a gene expression simulator
@@
    Input: gene expression data at four equally spaced time points
@@ 5 Performance of different methods
The same as in the WENDY paper and the TRENDY paper [9, 10], we test different methods on two synthetic data sets and two experimental data sets. SINC data set [10] and DREAM4 data set [44] are generated by simulating stochastic differential equation systems. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [45]. hESC data set is from human embryonic stem cell-derived progenitor cells[46]. THP-1 and hESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
@@
We test the performance of QWENDY, TEQWENDY, and LEQWENDY methods on these four data sets. Since TRENDY ranks the first among 16 methods on these four data sets [10], we also present the performance of TRENDY. For each method, we compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores [9]. These two scores, both between 00 and $1$, evaluate the level of matching, where $1$ means perfect match, and 00 means perfect mismatch.
@@ 6 Discussion
In this paper, we present QWENDY, a GRN inference method that requires single-cell gene expression data measured at four time points. Then we train the transformer-based variant TEQWENDY from scratch and fine-tune the LLM-based variant LEQWENDY based on pre-trained weights. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets. Notice that each experimental data set only has one group of data, meaning that the performance of each method has a large uncertainty level.
@@ Acknowledgments
For all models in TEQWENDY and LEQWENDY, the loss function is mean squared error; the optimizer is Adam with learning rate $0.001$; the number of training epochs is $100$. After each epoch, we evaluate the model performance on a validation set with 1000 samples. The training stops early and rolls back to the best status if there is no improvement for consecutive $10$ epochs.
@@
    Linear embedding layer with dimension $d$ to $d$ ($(n^{2})\times d$)
@@
    LeakyReLU activation function with $\alpha=0.1$ ($(n^{2})\times d$)
@@
    Linear embedding layer with dimension $d$ to $1$ ($(n^{2})\times 1$)